OpenStudy (anonymous):

HELP?! Input in standard form the equation of the given line. The line with m = -3/8 and passing through (5,-4)

OpenStudy (anonymous):

(X2-X1)(m)=(Y2-Y1)

OpenStudy (anonymous):

(5-x)(-3/8)=(-4-y)

OpenStudy (anonymous):

can u rearrange that into standard form now

OpenStudy (anonymous):

umm not really... but i know that the format is y = x + m, right?

OpenStudy (anonymous):

standard is in the format Ax+By+C no?

OpenStudy (mathmale):

Jessica: Why not look up and type out what you mean by "standard form" of the equation of a straight line? I'm used to the following: point-slope equation of a straight line, slope-intercept equation, general form (ax + by + c = 0). Once others and I know for certain which form you want, through what you type, it'd be easy for others to respond with helpful info.

OpenStudy (mathmale):

By the way: if I were given the slope m and one point on the line in question, I'd immediately substitute that info into the point-slope form.

OpenStudy (anonymous):

buy i don't know how to do with this equation m = -3/8 and passing through (5, -4)

OpenStudy (anonymous):

(-3/8) is your m and -4 would be ur b value

OpenStudy (mathmale):

Your most recent statement is correct, when m = -3/8 and b = -2. You used the "slope-intercept form." For the case where m = -3/8 and the line passes through (-5,-4), the proper form to use is "point slope." That's y-b = m(x-a), where (a,b) is the given point and m is the given slope. -4 is not the b value. Try again, please. m = -3/8; (a,b) = (-5,-4), and the correct form is y-b=m(x-a).

OpenStudy (anonymous):

i've been on this website answering quest for too long that my brain has malfunctioned i had the process right b4 with (5-x)(-3/8)=(-4-y) just need to expand that and rearrange

OpenStudy (mathmale):

For clarification: I've used (a,b) to represent the given point. Here that given point is (-5,-4), so (ngott3), -4 is the y-coordinate of the given point.

OpenStudy (anonymous):

First part of the equation answer is -3x - 8y right?

OpenStudy (mathmale):

I'd suggest you type out the whole thing so it's clear to everyone involved. The point-slope form of the equation of the line in question is y-(-4) = (-3/8)(x-(-5)). Please simplify that.

OpenStudy (anonymous):

-3x - 8y = -32

OpenStudy (anonymous):

omg i placed it the x and y values in the wrong spot

OpenStudy (anonymous):

yea i know my answer is wrong -.-"

OpenStudy (anonymous):

so what's the correct answer? I'm super confused...

OpenStudy (mathmale):

Jessica; y-(-4) = (-3/8)(x-(-5)) simplifies to y+4 = (-3/8)(x+5). Let's eliminate the fraction by multiplying both sides of the equation by 8: 8y + 32 = -3(x+5), or 8y + 32 = -3x -15 Re-arranging all terms so that they're on the left and we have 0 on the right, 3x + 8y + 47 = 0. This is in what I would call "general form". Don't take my word for its accuracy. Instead, check this equation by substituting the point (-5, -4) into 3x + 8y + 47 = 0 and determine whether or not the resulting equation is true. If it is, the equation is correct.

OpenStudy (anonymous):

yes, you are correct. The answer came out to 0

OpenStudy (anonymous):

Thanks so much!!!!!

OpenStudy (anonymous):

but how would i put that into standard form? @mathmale

OpenStudy (mathmale):

Jessica: I took my own advice and did an Internet search for "equation of a line in standard form," and ended up with the following quote: "Standard Form: the standard form of a line is in the form Ax + By = C where A is a positive integer, and B, and C are integers." Therefore, take my "3x + 8y + 47 = 0" and re-write it as 3x + 8y = -47. We both could have saved time by looking up and/or verifying that definition before doing anything else.